# Scizor: A Self-Supervised Approach to Data Curation for Large-Scale Imitation Learning

### v1

@@ Abstract
Imitation learning advances robot capabilities by enabling the acquisition of diverse behaviors from human demonstrations. However, large-scale datasets used for policy training often introduce substantial variability in quality, which can negatively impact performance. As a result, automatically curating datasets by filtering low-quality samples to improve quality becomes essential. Existing robotic curation approaches rely on costly manual annotations and perform curation at a coarse granularity, such as the dataset or trajectory level, failing to account for the quality of individual state-action pairs. To address this, we introduce Scizor, a self-supervised data curation framework that filters out low-quality state-action pairs to improve the performance of imitation learning policies. Scizor targets two complementary sources of low-quality data: suboptimal data, which hinders learning with undesirable actions, and redundant data, which dilutes training with repetitive patterns. Scizor leverages a self-supervised task progress predictor for suboptimal data to remove samples lacking task progression, and a deduplication module operating on joint state-action representation for samples with redundant patterns. Empirically, we show that Scizor enables imitation learning policies to achieve higher performance with less data, yielding an average improvement of $15.4$% across multiple benchmarks. More information is available at: https://ut-austin-rpl.github.io/SCIZOR/
@@ 1 Introduction
Imitation learning has shown promising signs in acquiring a wide range of motor behaviors by learning from expert demonstrations, a necessary step towards general-purpose robots. The success of the imitation learning hinges on the extensive datasets collected in a wide range of tasks and environments, often collected by different demonstrators. Such diverse, large-scale data collection inherently introduces variability in data quality [1, 2, 3], including mistakes made by operators leading to suboptimal actions (e.g., dropping an object), or redundancy in data leading to skewed distributions. Such a dataset can misguide models into learning incorrect behaviors [1, 4] and hinder diversity [3], reducing the impact of rare but informative actions. Therefore, effective data curation — the process of filtering data to improve the data quality [5, 6] — becomes critical for building robust and high-performing imitation learning policies.
@@
Figure 1: Scizor overview. Each trajectory from the original robotic datasets is simultaneously passed through the suboptimal transitions removal module and the redundant transitions removal module. Each module removes data based on its own threshold, resulting in a curated dataset. A policy trained on the curated dataset achieves a higher success rate.
@@
Early efforts in robotic data curation have relied heavily on human annotations to label high- and low-quality data [1], but these methods have largely been confined to small-scale datasets. As data scales up, manual annotation becomes infeasible, making it important to automatically curate data, an approach that has already shown promise in fields like computer vision (CV) and natural language processing (NLP) [7, 8, 9, 10]. Specifically in robot learning, a key challenge in this process to maximize data utilization is the need for curating at the finest granularity: we must evaluate the quality of individual state-action pairs. For instance, a trajectory may include an initial failed grasp attempt followed by a successful recovery, containing both suboptimal and valuable segments. Effective curation should isolate and remove only the uninformative or erroneous segments, like the failed grasp, while preserving segments that provide useful learning signals. The different impact of individual data points for learning has also been underscored in prior work, such as weighted behavior cloning [4, 11, 12]. However, current large-scale imitation learning curation methods have yet to address data quality at the transaction level. Existing approaches typically curate by reweighting entire dataset domains [13, 5] or discarding entire trajectories [6, 14], overlooking the contribution and quality of individual state-action pairs.
@@
Effectively curating individual state-action pairs is challenging, as robot demonstrations typically lack dense reward annotations, making it difficult to assess the quality of every interaction step. To address this, we adopt a self-supervised approach for filtering low-quality state-action pairs, offering a scalable solution for improving data quality in imitation learning. Our work is motivated by two key observations: (1) suboptimal transitions, which contain undesirable actions like collision, jittering, and other erroneous actions, can degrade policy performance by reinforcing incorrect behaviors; and (2) redundant transitions, which repeat common patterns excessively, can dilute the learning signal by dominating other informative and diverse samples.
@@
For effective large-scale imitation learning, we introduce Scizor, a self-supervised data curation method that reduces dataset size by filtering out suboptimal and redundant state-action pairs. First, to identify suboptimal data without access to reward information, we train a self-supervised task progress predictor using temporal distance classification [15, 16, 17], and remove frames that do not demonstrate meaningful progress toward the task goal. Second, to remove redundant data, a key insight is that some segments may appear visually similar while differing substantially in the executed actions. Therefore, both visual observations and their corresponding actions must be considered for effective deduplication. To this end, we apply deduplication [7] using joint representations of state and action to identify and filter redundant state-action pairs. We then filter out frames based on similarity scores to reduce repetition while preserving dataset diversity. The suboptimal frame filter targets harmful or noisy supervision, while the redundancy filter removes overrepresented patterns. Together, the two deletion strategies complement each other by targeting distinct modes of low-quality data.
@@
  • We propose a unified framework for data curation that filters both suboptimal and redundant state-action pairs in large-scale imitation learning datasets.
@@ 2 Related Work
Data quality has been known to affect policy learning performance for robotics [1, 43]. There have been studies in improving human demonstration quality, albeit in small-scale tasks, by automatic ranking [2], or eliciting compatible behavior from humans during the data collection process [44]. As progress in general-purpose, large-scale robot learning continues, there has been growing interest in curating large-scale datasets for robot learning [13, 23, 5, 6, 14]. Octo [13] and OpenVLA [23] perform ad-hoc dataset-level curation by heuristically tuning a set of weights for data mixtures, balancing the dataset composition; Remix [5] automates this dataset-level curation with distributionally robust optimization. DemInf [6] performs trajectory-level curation with mutual information as a trajectory quality estimator, and Demo-SCORE [14] also performs trajectory-level curation, but has to rely on online rollout performance.
@@ 3 Self-Supervised Data Curation for Large-Scale Imitation Learning
We introduce our data curation framework, Scizor, which performs fine-grained filtering of low-quality data in a self-supervised manner to improve imitation learning policy performance. We begin by introducing key formulations and background, followed by two core components of our method: (1) a self-supervised suboptimal transitions removal module and (2) a similarity-based state-action deduplication module that filters redundant transitions.
@@ 3.1 Preliminaries and Formulations
[…]
@@ 3.2 Suboptimal Transition Removal via Progress Estimation
Human demonstrations often contain both proficient and suboptimal segments in the same trajectory with no explicit signals. Manually labeling suboptimal segments is labor-intensive and not scalable. We propose a self-supervised approach to detect suboptimal behaviors based on the intuition that progress toward task completion should increase steadily over time. By training a model to predict task progress between pairs of observations, we learn a proxy for how much progress the agent makes over a given interval. If the predicted progress for a segment is unexpectedly lower than expected progress, it can serve as a signal of suboptimality. This allows us to automatically identify and filter out segments that deviate from making progress, without requiring any manual annotations.
@@
Defining Suboptimality with Task Progress. Inspired by temporal distance classification in self-supervised representation learning [15], we evaluate action quality by estimating the extent of task progress between two timesteps, $i$ and $i+T$. Specifically, we define a progress function $f:S_{i,i+T}\rightarrow T_{p}$, which inputs a sub-trajectory $S_{i,i+T}$ from timestep $i$ to $i+T$, and predicts the progress $T_{p}$ made over the sub-trajectory $S_{i,i+T}$ towards completion. Intuitively, $T_{p}$ measures the temporal distance that the robot has moved the task forward over the sub‑trajectory $S_{i,i+T}$, measured in seconds. We then compare this predicted progress $T_{p}$ to the actual elapsed time $T$. If the predicted progress is significantly lower than the elapsed time, this means the robot is behind schedule (i.e., progressing more slowly than expected), meaning that the sub-trajectory is suboptimal. The suboptimality score for the sub-trajectory $S_{i,i+T}$ is defined as $V_{i,i+T}=T-T_{p}$.
Predicting Task Progress. Rather than regressing a real-valued progress estimate, we cast progress prediction as classification over discrete temporal bins, which is empirically more robust [15]. We discretize the temporal gap into $B$ bins, where each bin is a time interval in seconds. To predict task progress for a sub-trajectory $S_{i,i+T}$, we train a task progress classifier to classify the bin corresponding to the time $T$ between the start and end states. Empirically, we set $B=5$, sample each sub-trajectory as $2$ seconds, and bins to be $[0,0.5),[0.5,1.0),[1.0,2.0),[2.0,5.0)$, and $[5.0,+\infty)$.
Assigning Suboptimal Scores to Individual Samples. Although we first compute a suboptimality score at the sub-trajectory level, our ultimate goal is to assign a suboptimality score to every individual transition to enable more fine-grained data filtering. Therefore, we aim to assign a sample-level suboptimal score $V_{i}$ to each transition at time $i$, based on the sub-trajectory-level suboptimal scores. We began by computing sub-trajectory-level suboptimal scores, $V_{0,T},V_{1,1+T}...,V_{N,N+T}$. Assuming that every sample within a sub-trajectory have equal probability to be suboptimal, each sub-trajectory score should be evenly distributed across its $T$ samples, for a sub-trajectory $S_{i,i+T}$ with $V_{i,i+T}$, it will contribute each transition within the sub-trajectory by $\frac{1}{T}V_{i,i+T}$. Therefore, the aggregated sample-level suboptimality score is then computed as: $\hat{V_{i}}=\sum_{t=i-T}^{i}\frac{1}{T}V_{t,t+T}$. Further, we apply the temporal discounting to our assigning, which gives us $V_{i}=\sum_{t=0}^{i}\gamma^{i-t}\hat{V_{t}}$ where $\gamma\in[0,1]$ is the discount factor. This captures the influence of past actions on the current state, while prioritizing more recent behavior.
Lastly, we observe that the overall quality of its trajectory can shape the quality of each action; for example, a human operator who is consistently expert-level or error-prone affects the quality of every action in the demonstration he/she collect. We therefore want each transition’s score to reflect not only its local quality but also the overall quality of the trajectory it belongs to. To compute the final curation score for each transition, we use a weighted combination of its suboptimality score with the mean score of the transitions across all timesteps, $\frac{1}{N}\sum_{j=1}^{N}(V_{j})$. We refer to this combination as the mixture of transition and trajectory scores. Specifically, the final score is computed as $0.5\times\frac{1}{N}\sum_{j=1}^{N}(V_{j})+0.5\times V_{i}$.
@@
Removing Suboptimal Samples. During policy training, we compute the suboptimality score for every sample as described above. During data curation, we define a suboptimal threshold $\epsilon_{s}$ and exclude transitions with suboptimality scores above this threshold from the training process. Note that each sample here preserves an observation history and an action sequence for algorithms that are history-dependent (e.g., BC-Transformer) and that utilize action chunking.
@@ 3.3 Similarity-Based State-Action Deduplication
A key insight is that some segments may appear visually similar, yet differ in task intent or executed actions. To avoid discarding meaningful variations, effective deduplication must consider both the visual states and actions. To this end, we propose a similarity-based deduplication method that utilizes joint representations of visual states and actions to identify and filter redundant state-action pairs.
Defining State-Action Duplicates. Prior work on semantic deduplication [7] has focused on curating large image datasets by removing semantically similar data pairs based solely on visual features. However, such visual-only deduplication methods are not well-suited for sequential decision-making tasks like imitation learning in robotics, where action dynamics play a crucial role. In this work, we extend the idea of semantic deduplication to the imitation learning domain by incorporating both visual states and action information. Specifically, we define state-action duplicates as state-action chunks $(S_{i,i+T},a_{i,i+T})$ that are visually similar and lead to comparable actions, reflecting redundant patterns that contribute little to learning diversity.
Generating State-Action Features. We first divide the dataset into non-overlapping sub-trajectories, each consisting of a state-action sequence $(S_{i,i+T},a_{i,i+T})$, where each chunk spans a fixed duration $T$. Given the variations in recording frequency across datasets, we uniformly subsample $N=8$ RGB images from each chunk for consistency. As raw visual data is high-dimensional and not directly suitable for similarity computation, we employ the Cosmos video encoder [45], a pre-trained model that encodes both temporal and semantic information from videos, to extract a compact 1D video feature vector $z_{v}$. We then concatenate the actions to the visual embedding to form a joint state-action feature $z_{v+a}$.
@@
3.4 Unified Threshold Across Datasets
When curating data, we remove all samples with a score above either $\epsilon_{s}$ or $\epsilon_{d}$. To generalize Scizor to different datasets, we need to find a unified threshold. We run Scizor on the RoboMimic and OXEMagic datasets under different curation thresholds and find that $\epsilon{s}=0.58$ and $\epsilon_{d}=0.99$ yield the best performance on both datasets. We then adopt this unified threshold for all experiments. For more details, please refer to Appendix A.

### v2

@@ Abstract
Imitation learning advances robot capabilities by enabling the acquisition of diverse behaviors from human demonstrations. However, large-scale datasets used for policy training often introduce substantial variability in quality, which can negatively impact performance. As a result, automatically curating datasets by filtering low-quality samples to improve quality becomes essential. Existing robotic curation approaches rely on costly manual annotations and perform curation at a coarse granularity, such as the dataset or trajectory level, failing to account for the quality of individual state-action pairs. To address this, we introduce Scizor, the first self-supervised transition-level curation framework that requires no annotations and scales to large-scale datasets to improve the performance of imitation learning policies and modern Vision-Language-Action (VLA) models. Scizor targets two complementary sources of low-quality data: suboptimal data, which hinders learning with undesirable actions, and redundant data, which dilutes training with repetitive patterns. Scizor leverages a self-supervised task progress predictor for suboptimal data to remove samples lacking task progression, and a deduplication module operating on joint state-action representation for samples with redundant patterns. Empirically, we show that Scizor enables imitation learning policies and modern VLA models to achieve higher performance with less data, yielding an average improvement of 15.415.4% across multiple benchmarks. More information is available at: https://ut-austin-rpl.github.io/SCIZOR/
@@ 1 Introduction
Imitation learning has shown promising signs in acquiring a wide range of motor behaviors by learning from expert demonstrations, a necessary step towards general-purpose robots. Recent advances in Vision-Language-Action (VLA) models have highlighted the potential of large-scale imitation learning, where massive datasets spanning diverse tasks and environments are used to train generalist policies. Such diverse, large-scale data collection inherently introduces variability in data quality [1, 2, 3], including mistakes made by operators leading to suboptimal actions (e.g., dropping an object), or redundancy in data leading to skewed distributions. Such a dataset can misguide models into learning incorrect behaviors [1, 4] and hinder diversity [3], reducing the impact of rare but informative actions. Therefore, effective data curation, the process of filtering data to improve the data quality [5, 6], becomes critical for building robust and high-performing imitation learning policies.
@@
Figure 1: Scizor overview. Each trajectory from the original robotic datasets is simultaneously passed through the suboptimal transitions removal module and the redundant transitions removal module. Each module removes data, resulting in a curated dataset. A policy trained on the curated dataset achieves a higher success rate.
@@
Early efforts in robotic data curation have relied heavily on human annotations to label high- and low-quality data [1], but these methods have largely been confined to small-scale datasets. As data scales up, manual annotation becomes infeasible, making it important to automatically curate data, an approach that has already shown promise in fields like computer vision (CV) and natural language processing (NLP) [7, 8, 9, 10]. Specifically in robot learning, a key challenge in this process to maximize data utilization is the need for curating at the finest granularity: we must evaluate the quality of individual state-action pairs. For instance, a trajectory may include an initial failed grasp attempt followed by a successful recovery, containing both suboptimal and valuable segments. Effective curation should isolate and remove only the uninformative or erroneous segments, like the failed grasp, while preserving segments that provide useful learning signals. The different impact of individual data points for learning has also been underscored in prior work, such as weighted behavior cloning [4, 11, 12]. However, current large-scale imitation learning curation methods have yet to address data quality at a fine-grained level. Existing approaches typically curate by reweighting entire dataset domains [13, 5] or discarding entire trajectories [6, 14], overlooking the contribution and quality of individual state-action pairs.
@@
Effectively curating individual state-action pairs is challenging, especially in large-scale datasets, as robot demonstrations typically lack dense reward annotations, making it difficult to assess the quality of every interaction step. To address this, we develop a self-supervised approach for filtering low-quality state-action pairs, offering a scalable solution for improving data quality in imitation learning. Our work is motivated by two key observations: (1) suboptimal transitions, which contain undesirable actions like collision, jittering, and other erroneous actions, can degrade policy performance by reinforcing incorrect behaviors; and (2) redundant transitions, which repeat common patterns excessively, can dilute the learning signal by dominating other informative and diverse samples.
@@
We introduce Scizor, the first self-supervised, transition-level data curation method that scales to the Open-X Embodiment Dataset [15] with more than one million trajectories for large-scale models such as the Octo [13] vision-language-action (VLA) model of over 27M parameters. Scizor is a self-supervised data curation method that reduces dataset size by filtering out suboptimal and redundant state-action pairs. First, to identify suboptimal data without access to reward information, we train a self-supervised task progress predictor using temporal distance classification [16, 17, 18], and remove frames that do not demonstrate meaningful progress toward the task goal. For instance, imagine a robot attempting to grasp an object and failing, then immediately trying again and succeeding. The initial failed attempt produces little or even negative progress, so it is discarded. The subsequent successful grasp, however, still advances the task and is retained. Second, to remove redundant data, a key insight is that some segments may appear visually similar while differing substantially in the executed actions, and vice versa. Therefore, both visual observations and their corresponding actions must be considered for effective deduplication. To this end, we apply deduplication [7] using joint representations of state and action to identify and filter redundant state-action pairs. We then filter out frames based on similarity scores to reduce repetition while preserving dataset diversity. In summary, the suboptimal frame filter targets harmful or noisy supervision, while the redundancy filter removes overrepresented patterns. Together, the two deletion strategies complement each other by targeting distinct modes of low-quality data.
@@
  • We propose a unified framework for transition-level data curation that filters both suboptimal and redundant state-action pairs in large-scale imitation learning datasets for imitation learning policies and modern Vision-Language-Action models.
@@ 2 Related Work
Data quality has been known to affect policy learning performance for robotics [1, 43]. There have been studies in improving human demonstration quality, albeit in small-scale tasks, by automatic ranking [2], or eliciting compatible behavior from humans during the data collection process [44]. As progress in general-purpose, large-scale robot learning continues, there has been growing interest in curating large-scale datasets for robot learning [13, 24, 5, 6, 14]. Octo [13] and OpenVLA [24] perform ad-hoc dataset-level curation by heuristically tuning a set of weights for data mixtures, balancing the dataset composition; Remix [5] automates this dataset-level curation with distributionally robust optimization. DemInf [6] performs trajectory-level curation with mutual information as a trajectory quality estimator, and Demo-SCORE [14] also performs trajectory-level curation, but has to rely on online rollout performance. DataMIL [45] select task relevant trajectories from Open-X [15] datasets to boost performance on task-specific policies.
@@ 3 Self-Supervised Data Curation for Large-Scale Imitation Learning
We introduce our data curation framework, Scizor, which performs fine-grained filtering of low-quality data in a self-supervised manner to improve imitation learning policy performance. We begin by introducing key formulations and background, followed by two core components of our method: (1) a self-supervised suboptimal transitions removal module; (2) a similarity-based state-action deduplication module that filters redundant transitions.
@@ 3.1 Preliminaries and Formulations
We formulate a robot manipulation task as a Markov Decision Process $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{R},\mathcal{P},p_{0},\gamma)$ representing the state space, action space, reward function, transition probability, initial state distribution, and discount factor. Given the current state $s_{t}\in\mathcal{S}$, the robot action $a_{t}\in\mathcal{A}$ is drawn from the policy $\pi\left(\cdot\mid s_{t}\right)$. The objective of imitation learning is to learn a policy $\pi_{R}$ parameterized by $\theta$ that maximizes the log-likelihood of actions $a$ conditioned on the states $s$:
$\theta^{*}=\underset{\theta}{\arg\max}\underset{(s,a)\sim\mathcal{D}_{expert}}{\mathbb{E}}\left[\log\pi_{\theta}(a\mid s)\right]$ (1)
where $(s,a)$ are samples from the human demonstration dataset $\mathcal{D}_{expert}$. Our data curation objective is to refine $\mathcal{D}_{expert}$ by filtering out suboptimal or redundant samples to improve policy performance.
@@ 3.2 Suboptimal Transition Removal via Progress Estimation
Human demonstrations often contain both proficient and suboptimal segments in the same trajectory with no explicit signals. Manually labeling suboptimal segments is labor-intensive and not scalable. We propose a self-supervised approach to detect suboptimal behaviors based on the intuition that progress toward task completion should increase steadily over time. As we don’t have ground-truth progress labels, we use the ground truth time elapsed between two observations as a stand-in for the task progress. We train a lightweight model to predict this progress from pairs of states. At test time, if the predicted progress for a segment is unexpectedly lower than expected progress, it can serve as a signal of suboptimality. This allows us to automatically identify and filter out segments that deviate from making progress, without requiring any manual annotations.
@@
Defining Suboptimality with Task Progress. Inspired by temporal distance classification in self-supervised representation learning [16], we evaluate action quality by estimating the task progress between two timesteps. If the predicted progress is significantly lower than the elapsed time, this means the robot is behind schedule (i.e., progressing less than expected), meaning that the sub-trajectory is suboptimal. Specifically, we define a progress function $f:S_{i:i+T}\rightarrow T_{p}$, which inputs a sub-trajectory $S_{i:i+T}$ from timestep $i$ to $i+T$, and predicts the progress $T_{p}$ made over the sub-trajectory $S_{i:i+T}$ towards completion. $T_{p}$ measures the temporal distance that the robot has moved the task forward over the sub‑trajectory $S_{i:i+T}$, measured in seconds. We then compare this predicted progress $T_{p}$ to the actual elapsed time $T$. The suboptimality score for the sub-trajectory $S_{i:i+T}$ is defined as $V_{i:i+T}=T-T_{p}$.
Assigning Suboptimal Scores to Individual Samples. Our goal is to assign a suboptimality score to each individual transition, enabling fine-grained data filtering. While our scores are initially predicted at the sub-trajectory level, we want each transition’s score to include 3 factors: (1) Progress predicted from each sub-trajectory that current transition belongs to, (2) Future suboptimality with reducing impact over time, so that the current transition score accounts for suboptimal scenarios it leads to, (3) The overall quality of the entire trajectory that the transition belongs to. This enables a transition’s score to reflect both its local impact and the global quality of the whole trajectory.
Formally, from the previous section we have sub-trajectory scores $V_{0:T},V_{1:1+T},\dots,V_{N:N+T}$. Assuming equal contribution from each transition in a sub-trajectory, a score $V_{i:i+T}$ is evenly distributed across its $T$ constituent transitions, contributing $\frac{1}{T}V_{i:i+T}$ to each. The aggregated sample-level score is computed as $\hat{V_{i}}=\sum_{t=i-T}^{i}\frac{1}{T}V_{t:t+T}.$ To allow the current suboptimality score to also account for future suboptimality with diminishing impact over time, we apply temporal discounting: $V_{i}=\sum_{t=i}^{T}\gamma^{t-i}\hat{V_{t}},\quad\gamma\in[0,1],$ where $\gamma$ controls the rate at which future influences diminish. Finally, to incorporate trajectory-level quality, we combine each transition’s score $V_{i}$ with the mean score across all $N$ transitions in its trajectory, yielding the final curation score:
$V_{i}^{\text{final}}=\alpha\cdot V_{i}+(1-\alpha)\cdot\frac{1}{N}\sum_{j=1}^{N}V_{j},$
where $\alpha\in[0,1]$ balances local and trajectory-wide quality.
@@
Removing Suboptimal Samples. During policy training, we compute the suboptimality score for every sample as described above. During data curation, we exclude transitions with suboptimality scores above a certain threshold $\epsilon_{s}$ from the training process.
@@ 3.3 Similarity-Based State-Action Deduplication
A key insight is that some segments may appear visually similar, yet differ in task intent or executed actions. To avoid discarding meaningful variations, effective deduplication must consider both visual states and actions. To this end, we propose a similarity-based deduplication method that utilizes joint representations of visual states and actions to identify and filter redundant state-action pairs.
Defining State-Action Duplicates. Prior work on semantic deduplication [7] has focused on curating large image datasets by removing semantically similar data pairs based solely on visual features. However, such visual-only deduplication methods are not well-suited for sequential decision-making tasks like imitation learning in robotics, where action dynamics play a crucial role. In this work, we extend the idea of semantic deduplication to the imitation learning domain by incorporating both visual states and action information. Specifically, we define state-action duplicates as state-action chunks $(S_{i:i+T},a_{i:i+T})$ that are visually similar and lead to comparable actions, reflecting redundant patterns that contribute little to learning diversity.
Generating State-Action Features. We first divide the dataset into non-overlapping sub-trajectories, each consisting of a state-action sequence $(S_{i:i+T},a_{i:i+T})$, where each chunk spans a fixed duration $T$. Given the variations in recording frequency across datasets, we uniformly subsample $N$ RGB images from each chunk for consistency. As raw visual data is high-dimensional and not directly suitable for similarity computation, we employ the Cosmos video encoder [46], a pre-trained model that encodes both temporal and semantic information from videos, to extract a compact 1D video feature vector $z_{v}$. We then concatenate the actions represented as delta end-effector pose to the visual embedding to form a joint state-action feature $z_{v+a}$.